# Metalic: Meta-Learning In-Context with Protein Language Models

### v1

@@ 1 Introduction
The accurate prediction of functional and biophysical properties of proteins, such as stability or binding affinity, is a critical challenge in the physical sciences with far-reaching implications for drug discovery, medical research, and agriculture. Protein design seeks to optimize one, or a collection of these properties, collectively referred to as fitness. While protein fitness can be measured in vitro, the process is laborious and time-consuming. Consequently, machine learning models have emerged as a powerful tool to accelerate protein design by predicting fitness directly from amino acid sequences in silico. However, due to the complex, high-dimensional relationship between protein sequences and fitness, and the limited availability of high-quality data, accurate fitness prediction is a challenge.
@@
Proteins can be encoded as sequences of characters representing amino acids, making protein language models (PLMs) effective for modeling them (Madani et al., 2020; Rives et al., 2021; Rao et al., 2021; Lin et al., 2022; Notin et al., 2023; Truong Jr & Bepler, 2024). By predicting masked amino acids, or subsequent amino acids, over known proteins at scale, PLMs can capture much of the structure and resultant properties deriving from the amino acid sequence. While PLMs are not directly trained to predict fitness, they are trained to model the likelihood of naturally occurring proteins, which has been found to correlate strongly with their fitness (Truong Jr & Bepler, 2024). In practice, large PLMs are fine-tuned on downstream protein fitness data, often with linear regression heads added (Krause et al., 2022). While this paradigm is highly effective, when there is severely limited data, or no data, PLMs provide limited utility. With limited data, learning the correct regression is difficult, even with informative pre-trained representations. With no data, it must be assumed that the protein fitness is solely a function of protein likelihood (Truong Jr & Bepler, 2024; Hawkins-Hooker et al., 2024), and for masked language models, it is also assumed that each amino acid contributes independently to fitness (Meier et al., 2021; Notin et al., 2024).
@@
While limited or no data may be available for specific protein fitness prediction tasks, there are often other tasks that can serve as a valuable source of guidance. For example, rather than assuming that fitness is solely a function of protein likelihood, when no data is available for a given task, we can use data from other tasks to learn the relationship between PLM embeddings and fitness. Due to advances in high-throughput assays, such as deep mutational scanning Fowler & Fields (2014), data from other tasks is available to learn this relationship. One example is ProteinGym (Notin et al., 2024), which compiles over one hundred distinct fitness prediction tasks. Training over such a distribution of related tasks, to reason about new tasks, is referred to as meta-learning (Huisman et al., 2021; Hospedales et al., 2021; Beck et al., 2023).
@@ 2 Related Work
Gradient-based algorithms and in-context algorithms differ both in their computational efficiency and capacity for out-of-distribution generalization. Gradient-based approaches explicitly adapt model parameters within in the inner loop using standard gradient-based learning. Commonly, the model initialization are the only parameters learned over the task distribution. The meta-training process explicitly computes gradients through this adaptation process, effectively learning a parameter initialization that can be adapted to new tasks with only a few gradient steps (Finn et al., 2017; Zintgraf et al., 2019; Vuorio et al., 2019). However, this comes with considerable computational overhead – in particular when calculating meta-gradients through the inner loop – which makes gradient-based meta-learning less suitable for large models. Alternatively, in-context meta-learning adapts by conditioning on a task-specific dataset in context, i.e., conditioning on the data points over which gradient-based approaches would train. Typically, this is achieved by modelling the inner loop with a sequence model that can explicitly condition on the provided data (Santoro et al., 2016; Mishra et al., 2017; Beck et al., 2024a; b). Such methods are typically more sample- and compute-efficient than gradient-based methods, but can perform worse when provided with out-of-distribution tasks given the lack of explicit gradient-based learning in the inner-loop (Beck et al., 2023).
@@ 3.2 Metalic
Architecture Our work leverages the ProteinNPT architecture proposed by Notin et al. (2023) as an in-context PLM for fitness prediction. Here we briefly summarise the key elements, but defer the reader to Section A.2 and the original paper for full details. The architecture, along with the data we use for meta-leaning, are illustrated in Fig. 2(a).
Protein sequences in both the support and query set are converted to per-residue embeddings using a pre-trained PLM; in our case we take the third layer of ESM2-8M (Lin et al., 2022). Fitness scores in the query set are set to zero. All fitness scores are then projected to match the dimension of the residue embedding with a linear layer, summed with a learned embedding denoting if the row belongs to the support or query set, and finally concatenated along the sequence dimension. Optionally, zero-shot fitness predictions from an auxiliary PLM can be embedded and concatenated in the same way, which we explore in Section 4.3. This tensor is then processed via axial attention blocks (Ho et al., 2019), each of which applies self-attention separately along and across the sequences. Axial attention reduces the computational complexity of self-attention from to , where is the shot and is the length of a protein. Finally, a multi-layer perceptron conditions on the fitness embedding and mean-pooled sequence embedding to predict each query value, i.e. , which is then used to rank the proteins by fitness.
@@ 4.5 Ablations
We also ablate the ability to attend to the rest of the proteins in context by turning off the column attention in the axial attention layers (NoICL), we ablate the fine-tuning stage of training (NoFT), we ablate the preference-based loss by using mean squared error instead (NoPref), and we ablate the augmentation of single-mutant training data with multi-mutant data (NoAug). Note that when we remove the fine-tuning, and have a non-zero support size, we allow the early stopping data to be passed in-context, to not unfairly advantage fine-tuning with additional data. Ablating in-context learning decreases performance in both settings, indicating an ability to adapt in an unsupervised fashion even from the query set alone. Ablating fine-tuning decreased performance in the 128-shot setting, indicating the need for fine-tuning for generalization to held-out data. Ablating the preference-based loss decreased performance in both settings, in line with recent literature (Krause et al., 2022; Brookes et al., 2023; Hawkins-Hooker et al., 2024). Finally, ablating the multi-mutant augmentation decreased performance in the zero-shot setting, with comparable 128-shot performance, indicating the benefit of additional meta-training data when fine-tuning data is most limited.
@@ 4.6 Gradient-Based Meta-Learning
| Metalic (45k) | 45,000 | 45,000 | .553 .003 |
| Metalic (15k) | 15,000 | 15,000 | .512 .003 |
| Reptile-3-3 | 15,000 | 45,000 | .504 .009 |
| Reptile-3-100 | 15,000 | 45,000 | .528 .004 |
| Metalic-Reptile | 15,000 | 45,000 | .532 .006 |
Although Reptile is more compute-efficient than other gradient-based methods, it still performs gradient updates in the inner loop during meta-training, making it computationally expensive relative to Metalic. Our method uses 100 updates on the support data for fine-tuning. Reptile uses inner-loop gradient updates during both meta-training and fine-tuning. Due to compute limitations, we cannot use 100 steps for each forward pass of meta-training. Thus, we evaluate Reptile with 3 inner-loop gradient steps during meta-training and 3 during fine-tuning, so that train and test match (Reptile-3-3), and we evaluate Reptile with 3 inner-loop gradient steps during meta-training and 100 during fine-tuning, so that test time matches our method (Reptile-3-100). Note that even Reptile-3-3 uses three times more compute than Metalic. Finally, we evaluate Reptile-3-100 with Metalic, to see whether they can be used in conjunction (Metalic-Reptile). We train all methods for 15,000 steps, given the increased computation, and also report Metalic after 45,000 steps, to match the same number of gradient computations as the Reptile models.
Results are reported in Table 6. We evaluate in the 128-shot setting. Note that Reptile is not applicable in the zero-shot setting, as it requires some data for fine-tuning. We observe that without controlling for the amount of computation, Metalic (15k) performs similarly to Reptile, with a Spearman correlation between Reptile-3-3 and Reptile-3-100. We also observe that Metalic, in conjunction with Reptile (Metalic-Reptile), outperforms Reptile and Metalic independently. However, the improvement over Metalic is marginal, and the increased computation (by a factor of 3) is large. Controlling for the total number of gradient computations, Metalic (45k) achieves a Spearman correlation much higher than the combined Metalic-Reptile. Thus, while Reptile can be used in conjunction with Metalic, it may not be worth the added computation. This result is in line with previous work showing that accounting for gradient updates during meta-training (in their case, without in-context learning) can be detrimental when there are a limited number of tasks for meta-training or a limited amount of data for the inner-loop (Gao & Sener, 2020; Triantafillou et al., 2020). Collectively, this evidence further justifies fine-tuning only after meta-training.
@@ 6 Conclusion
In this paper, we have demonstrated state-of-the-art results on a standard protein fitness prediction benchmark in low-data settings. In order to do so, we proposed Metalic, which makes use of both in-context meta-learning and subsequent fine-tuning. Critically, we have demonstrated the ability of meta-learning to take advantage of additional data from other proteins fitness prediction tasks, while remaining computationally tractable by deferring fine-tuning to test time alone. Unique within the meta-learning literature, we show that in-context meta-learning provides a useful initialization for further fine-tuning, and can make use of test time data for both fine-tuning and in-context learning. Metalic additionally demonstrates the ability to learn from the query set alone (zero-shot), performing unsupervised in-context learning. Future work could investigate leveraging Metalic’s unique in-context learning ability to act as an auto-regressive fitness model (i.e., a world model in the meta-reinforcement learning setting (Beck et al., 2023)), for optimally trading off exploration and exploitation when designing novel proteins. Given its efficacy at leveraging additional data, we believe that meta-learning in-context will play a crucial role in advancing protein fitness prediction, with Metalic being a foundational step in that direction.
@@ A.2 Hyper-Parameters
Since we build upon the axial attention of ProteinNPT (Notin et al., 2023), we follow their choice for most hyper-parameters, with a few exceptions. Most notably, in our experiments, we use the third layer of ESM2 as an embedding for each protein, given the strong performance and reduced number of parameters. Additionally, we use a ranking loss from Hawkins-Hooker et al. (2024), and do not use the CNN or additional inputs (such as zero-shot predictions) from ProteinNPT. We likewise found conditioning on the wild-type unhelpful. The same set of hyper-parameters are used for each setting: 0-shot, 16-shot, and 128-shot, and for multi-mutant results. We used the same learning rate for fine-tuning Metalic as for meta-training. We found the default ProteinNPT learning rate to be too large, and decreasing by a factor of 5 to be sufficient. Complete details on hyper-parameters used are available in Table 7. We tuned relatively few of the hyper-parameters of our method, and mostly tuned over a single seed. There is likely room for improvement in the hyper-parameter selection of Metalic.
@@
For the majority of baselines no tuning was required, other than Reptile. For the baselines, we used reference predictions for Table 1 and reference implementations for Table 2, neither of which required tuning. For Reptile, the update in the outer-loop can be uses as written in Equation 5, or can be interpreted as a gradient for use with the Adam optimizer Nichol et al. (2018). We experiment with both and find the use of the Adam optimizer to be superior in performance. All results in the main body use the Adam optimizer for Reptile. In order to evaluate the method without the Adam optimizer, we re-tune in Equation 5, which is the outer-loop learning rate. We leave the inner-loop learning rate, , fixed at the learning rate for Metalic, , which uses the same learning rate for the inner- and outer-loop. Given the increased computational cost of Reptile, we use a single seed over three learning rates for 10,000 steps. We tune over the following learning rates (): the learning rate for Metalic, which is ; a learning rate of 1, which corresponds to the learning rate of Metalic if you interpret as the gradient in a gradient-based update (Nichol et al., 2018); and a learning rate of , where is the number of inner-loop updates (3 in this case), which corresponds to the learning rate of Metalic if you interpret as the gradient in a gradient-based update111he standard gradient update, , with , gives , which is the same update as Equation 5 with , omitting the expectation for brevity.. We found a learning rate of 1 to perform best, in line with the outer loop learning rate of Metalic and the interpretation of the gradient from Nichol et al. (2018). For Adam, there is no , and we do not re-tune Adam’s outer-loop learning rate. When using Adam, the gradient is interpreted as . The results of tuning the learning rate without Adam, suggests a learning rate of with this gradient interpretation. Indeed, we confirm that when using Adam with this learning rate, and this gradient interpretation, it works better than using the alternative Reptile update rule with any learning rate. Results of the optimizer and learning rate tuning are presented in table Table 8. Note that these results used an earlier set of hyper-parameters.
@@
A.3 Fine-Tuning Warm-Up
@@
A.4 More Multi-Mutant Results

### v2

@@ 1 Introduction
The accurate prediction of functional and biophysical properties of proteins, collectively referred to here as fitness, is a critical challenge in the physical sciences with far-reaching implications for medical research, agriculture, and drug discovery. For example, fitness prediction can be used to optimize properties such as the binding affinity of a monoclonal antibody therapy to its target or the thermostability of enzymes functioning at high temperatures. While protein fitness can be measured in vitro, the process is laborious and time-consuming. Consequently, machine learning models have emerged as a powerful tool to predict fitness directly from amino acid sequences in silico. However, due to the complex, high-dimensional relationship between protein sequences and fitness, and the limited availability of high-quality data, accurate fitness prediction is a challenge.
@@
Proteins can be encoded as sequences of characters representing amino acids, making protein language models (PLMs) effective for modeling them (Madani et al., 2020; Rives et al., 2021; Rao et al., 2021; Lin et al., 2022; Notin et al., 2023; Truong Jr & Bepler, 2024). By predicting masked amino acids, or subsequent amino acids, over known proteins at scale, PLMs can capture much of the structure and resultant properties deriving from the amino acid sequence. While PLMs are not directly trained to predict fitness, they are trained to model the likelihood of naturally occurring proteins, which has been found to correlate strongly with their fitness (Truong Jr & Bepler, 2024). In practice, large PLMs are fine-tuned on downstream protein fitness data for regression. While highly effective, PLMs provide limited utility given severely limited data, or no data. With limited data, learning the correct regression is difficult, even with informative pre-trained representations. With no data, it must be assumed that protein fitness is solely a function of protein likelihood (Truong Jr & Bepler, 2024; Hawkins-Hooker et al., 2024), and for masked language models, it is also assumed that each amino acid contributes independently to fitness (Meier et al., 2021; Notin et al., 2024).
@@
While limited or no data may be available for specific protein fitness prediction tasks, there are often other tasks that can serve as a valuable source of guidance. For example, rather than assuming that fitness is solely a function of protein likelihood, when no data is available for a given task, we can use data from other tasks to learn the relationship between PLM embeddings and fitness. Due to advances in high-throughput assays, such as deep mutational scanning Fowler & Fields (2014), data from other tasks is available to learn this relationship. One example is ProteinGym (Notin et al., 2024), which compiles over one hundred distinct fitness prediction tasks. Training over such a distribution of related tasks, to reason about new tasks, is referred to as meta-learning (Huisman et al., 2021; Hospedales et al., 2021; Beck et al., 2023a).
@@ 2 Related Work
Gradient-based and in-context algorithms differ both in their computational efficiency and capacity for out-of-distribution generalization. Gradient-based approaches explicitly adapt model parameters within in the inner loop using standard gradient-based learning. Commonly, a parameter initialization is learned that can be adapted to new tasks with only a few gradient steps (Finn et al., 2017; Zintgraf et al., 2019; Vuorio et al., 2019). However, this comes with considerable computational overhead – due to meta-gradients when differentiating the inner loop – which makes gradient-based meta-learning less suitable for large models. Alternatively, in-context meta-learning adapts by conditioning a sequence model on a task-specific dataset in context. These methods condition on the data points over which gradient-based approaches would train (Santoro et al., 2016; Mishra et al., 2017; Beck et al., 2024; 2023b). Such methods are typically more sample- and compute-efficient than gradient-based methods, but perform worse on out-of-distribution tasks given the lack of explicit gradient-based learning in the inner loop (Beck et al., 2023a). Rather than training for in-context learning, the in-context learning of pre-trained large language models can also be used to perform meta-learning (Coda-Forno et al., 2023); however, this lacks generalization guarantees and requires fitting all tasks in context simultaneously, which is not possible given our data (Section 4.1).
@@ 3.2 Metalic
Architecture Our work leverages the ProteinNPT architecture proposed by Notin et al. (2023) as an in-context PLM for fitness prediction. Here we briefly summarize the key elements, but defer the reader to Section A.2 and the original paper for full details. The architecture, along with the data we use for meta-leaning, are illustrated in Fig. 2(a). A detailed illustration is give in Section A.6.
Protein sequences in both the support and query set are converted to per-residue embeddings using a pre-trained PLM; in our case we take the third layer of ESM2-8M (Lin et al., 2022). Fitness scores in the support set are projected to match the dimensionality of the residue embeddings using a linear layer, while the query set fitness embeddings share a single learned embedding. The protein sequence embeddings and fitness embeddings are then concatenated along the sequence dimension. Optionally, zero-shot fitness predictions from an auxiliary PLM can be embedded and concatenated in the same way, which we explore in Section 4.3. This tensor is then processed via axial attention blocks (Ho et al., 2019), each of which applies self-attention separately along and across the sequences. Axial attention reduces the computational complexity of self-attention from to , where is the shot and is the length of a protein. Finally, a multi-layer perceptron conditions on the fitness embedding and mean-pooled sequence embedding to predict each query value, i.e. , which is then used to rank the proteins by fitness.
@@ 4.5 Ablations
We also ablate the ability to attend to the rest of the proteins in context by turning off the column attention in the axial attention layers (NoICL), we ablate the fine-tuning stage of training (NoFT), and we ablate the preference loss by using mean squared error instead (NoPref). Note that when we remove the fine-tuning, and have a non-zero support size, we allow the early stopping data to be passed in-context, to not unfairly advantage fine-tuning with additional data. Ablating in-context learning decreases performance in both settings, indicating an ability to adapt in an unsupervised fashion even from the query set alone. Ablating fine-tuning decreased performance in the 128-shot setting, indicating the need for fine-tuning for generalization to held-out data. Ablating the preference-based loss decreased performance in both settings, in line with recent literature (Krause et al., 2022; Brookes et al., 2023; Hawkins-Hooker et al., 2024). An additional ablation is given in Section A.3.
@@ 4.6 Gradient-Based Meta-Learning
| Metalic (150k) | 150,000 | 150,000 | .562 .004 |
| Metalic-Reptile | 50,000 | 150,000 | .562 .004 |
| Metalic (50k) | 50,000 | 50,000 | .552 .009 |
| Reptile-3-100 | 50,000 | 150,000 | .539 .001 |
| Reptile-3-3 | 50,000 | 150,000 | .499 .008 |
Although Reptile is more compute-efficient than other gradient-based methods, it still performs gradient updates in the inner loop during meta-training, making it computationally expensive relative to Metalic. Our method uses 100 updates on the support data for fine-tuning. Reptile uses inner-loop gradient updates during both meta-training and fine-tuning. Due to compute limitations, we cannot use 100 steps for each forward pass of meta-training. Thus, we evaluate Reptile with 3 inner-loop gradient steps during meta-training and 3 during fine-tuning, so that train and test match (Reptile-3-3), and we evaluate Reptile with 3 inner-loop gradient steps during meta-training and 100 during fine-tuning, so that test time matches our method (Reptile-3-100). Note that even Reptile-3-3 uses three times more compute than Metalic. Finally, we evaluate Reptile-3-100 with Metalic, to see whether they can be used in conjunction (Metalic-Reptile). We train for 50,000 steps, and also report Metalic after 150,000 steps, to match the number of gradient computations as the Reptile models.
Results are reported in Table 6. We evaluate in the 128-shot setting. Note that Reptile is not applicable in the zero-shot setting, as it requires some data for fine-tuning. We observe that without controlling for the amount of computation, Metalic (50k) outperforms both Reptile variants. We also observe that Metalic, in conjunction with Reptile (Metalic-Reptile), outperforms Reptile and Metalic independently. However, the improvement over Metalic is marginal, and the increased computation (by a factor of 3) is large. Controlling for the total number of gradient computations, Metalic (150k) achieves a Spearman correlation comparable to the combined Metalic-Reptile. Thus, Reptile performs poorly in isolation, and, while Reptile can be used with Metalic, the combination increases implementation complexity with no clear advantage. This result is in line with previous work showing that accounting for gradient updates during meta-training (in their case, without in-context learning) can be detrimental when there are a limited number of tasks for meta-training or a limited amount of data for the inner-loop (Gao & Sener, 2020; Triantafillou et al., 2020). Collectively, this evidence further justifies fine-tuning only after meta-training.
@@ 6 Conclusion
In this paper, we have demonstrated state-of-the-art results on a standard protein fitness prediction benchmark in low-data settings. To do so, we proposed Metalic, which makes use of both in-context meta-learning and subsequent fine-tuning. Critically, we have demonstrated the ability of meta-learning to take advantage of additional data from other proteins fitness prediction tasks, while remaining computationally tractable by deferring fine-tuning to test time alone. Unique within the meta-learning literature, we show that in-context meta-learning provides a useful initialization for further fine-tuning, and can make use of test time data for both fine-tuning and in-context learning. Metalic additionally demonstrates the ability to learn from the query set alone (zero-shot), performing unsupervised in-context learning. Future work could investigate leveraging Metalic’s unique in-context learning ability to act as an auto-regressive fitness model (i.e., a world model in the meta-reinforcement learning setting (Beck et al., 2023a)), for optimally trading off exploration and exploitation when designing novel proteins. Given its efficacy at leveraging additional data, we believe that meta-learning in-context will play a crucial role in advancing protein fitness prediction, with Metalic being a foundational step in that direction.
@@ A.2 Hyper-Parameters
Since we build upon the axial attention of ProteinNPT (Notin et al., 2023), we follow their choice for most hyper-parameters, with a few exceptions. Most notably, in our experiments, we use the third layer of ESM2 as an embedding for each protein, given the strong performance and reduced number of parameters. Additionally, we use a ranking loss from Hawkins-Hooker et al. (2024), and do not use the CNN or additional inputs (such as zero-shot predictions) from ProteinNPT. We likewise found conditioning on the wild-type unhelpful. The same set of hyper-parameters are used for each setting: 0-shot, 16-shot, and 128-shot, and for multi-mutant results. We used the same learning rate for fine-tuning Metalic as for meta-training. We found the default ProteinNPT learning rate to be too large, and decreasing by a factor of 5 to be sufficient. Consistent with our baselines, we select hyper-parameters using the single-mutant results and then use the same hyper-parameters for the multi-mutant setting, following Hawkins-Hooker et al. (2024); Notin et al. (2023). Complete details on hyper-parameters used are available in Table 7. We tuned relatively few of the hyper-parameters of our method, and mostly tuned over a single seed. There is likely room for improvement in the hyper-parameter selection of Metalic.
@@
For the majority of baselines no tuning was required, other than Reptile. For the baselines, we used reference predictions for Table 1 and reference implementations for Table 2, neither of which required tuning. For Reptile, the update in the outer-loop can be uses as written in Equation 5, or can be interpreted as a gradient for use with the Adam optimizer Nichol et al. (2018). We experiment with both and find the use of the Adam optimizer to be superior in performance. All results in the main body use the Adam optimizer for Reptile. In order to evaluate the method without the Adam optimizer, we re-tune in Equation 5, which is the outer-loop learning rate. We leave the inner-loop learning rate, , fixed at the learning rate for Metalic, , which uses the same learning rate for the inner- and outer-loop. Given the increased computational cost of Reptile, we use a single seed over three learning rates for 10,000 steps. We tune over the following learning rates (): the learning rate for Metalic, which is ; a learning rate of 1, which corresponds to the learning rate of Metalic if you interpret as the gradient in a gradient-based update (Nichol et al., 2018); and a learning rate of , where is the number of inner-loop updates (3 in this case), which corresponds to the learning rate of Metalic if you interpret as the gradient in a gradient-based update111The standard gradient update, , with , gives , which is the same update as Equation 5 with , omitting the expectation for brevity.. We found a learning rate of 1 to perform best, in line with the outer loop learning rate of Metalic and the interpretation of the gradient from Nichol et al. (2018). For Adam, there is no , and we do not re-tune Adam’s outer-loop learning rate. When using Adam, the gradient is interpreted as . The results of tuning the learning rate without Adam, suggests a learning rate of with this gradient interpretation. Indeed, we confirm that when using Adam with this learning rate, and this gradient interpretation, it works better than using the alternative Reptile update rule with any learning rate. Results of the optimizer and learning rate tuning are presented in table Table 8. Note that these results used an earlier set of hyper-parameters.
@@
A.3 Augmentation Ablation
In this section we provide an additional ablation. Specifically, we ablate the augmentation of single-mutant training data with multi-mutant data (NoAug). The result is reported in Table 9. Ablating the multi-mutant augmentation decreased performance in the zero-shot setting, with comparable 128-shot performance, indicating the benefit of additional meta-training data when fine-tuning data is most limited.
| Model Name | ||
| Metalic | .482 .002 | .552 .009 |
| Metalic-NoAug | .464 .011 | .558 .002 |
A.4 Fine-Tuning Warm-Up
@@
A.5 More Multi-Mutant Results
@@
A.6 Detailed Architecture
Here, we present Fig. 6, which gives a more detailed version of the architecture diagram.